**THERE are 14 boys and 16 girls in Mr. Heath's class. The name of each student in the class was written on a slip of paper. Mr. Heath will draw a name from the bag, and WITHOUT replacing it, draw a second name. What is the probability that he will pick a boy followed by a girl?
a.1/124
b.112/435
c.2/3
d.435/443

Answer by sudhanshu_kmr(1152)


14+16 = 30
probability = (14/30) * (16/29) = 112/435
ans. b